# GEGA: Graph Convolutional Networks and Evidence Retrieval Guided Attention for Enhanced Document-level Relation Extraction

### v1

@@ Abstract
Yanxu Mao1, Peipei Liu2 ††thanks: Corresponding author., Tiehan Cui1 1School of Software, Henan University, China 2Zhongguancun Laboratory, Beijing, China {maoyanxu,cuitiehan}@henu.edu.cn peipliu@yeah.net
@@ 1 Introduction
(1) We propose a novel DocRE model, GEGA111The implementation code for GEGA can be obtained from the GitHub link: [Links covered up in blind reviewing] (Graph Convolutional Networks and Evidence Retrieval Guided Attention). This model combines graph structures and Transformers to retrieve evidence sentences highly relevant to the relational entity pairs from the document, guiding the attention to assign higher weights to this evidence information, thereby enhancing the performance of relation extraction.
@@ 4.1 Input Encoder Layer
For each entity, the efficacy of the max pooling function is pronounced when the inter-entity relations are explicitly articulated. Nevertheless, in the context of this study, the relations among entities remain ambiguous. It is understood that an entity may be referenced by one or several mentions, and a mention might uniquely identify an entity or fail to ascertain a definite corresponding entity. This necessitates the calculation of an entity’s embedding based on the embeddings of each associated mention. Following methodology of Jia et al. (2019), a soft version of the max function , is utilized to compute the embeddings of entities while concurrently capturing the attention distribution:
where is an entity comprising multiple mentions, while is the distribution of attention for mention within entity . is the number of mentions for entity .
@@ 4.2.4 Collaborative Prediction Layer
where is the row vector composed of all , is the total count of sentences contained within the document.

### v2

@@ Abstract
Yanxu Mao1, Xiaohui Chen2, Peipei Liu3,4 ††thanks: Corresponding author., Tiehan Cui1, Zuhui Yue2, Zheng Li2 1School of Software, Henan University, Kaifeng, China 2China Mobile Research Institute, Beijing, China 3Institute of Information Engineering, Chinese Academy of Sciences, Beijing, China 4School of Cyber Security, University of Chinese Academy of Sciences, Beijing, China {maoyanxu,cuitiehan}@henu.edu.cn xiaohuichen1116@gmail.com peipliu@yeah.net {yuezuhui, lizheng}@chinamobile.com
@@ 1 Introduction
(1) We propose a novel DocRE model, GEGA111The implementation code for GEGA can be obtained from the GitHub link: : https://github.com/maoxuxu/GEGA (Graph Convolutional Networks and Evidence Retrieval Guided Attention). This model combines graph structures and Transformers to retrieve evidence sentences highly relevant to the relational entity pairs from the document, guiding the attention to assign higher weights to this evidence information, thereby enhancing the performance of relation extraction.
@@ 4.1 Input Encoder Layer
For each entity, the efficacy of the max pooling function is pronounced when the inter-entity relations are explicitly articulated. Nevertheless, in the context of this study, the relations among entities remain ambiguous. It is understood that an entity may be referenced by one or several mentions, and a mention might uniquely identify an entity or fail to ascertain a definite corresponding entity. This necessitates the calculation of an entity’s embedding based on the embeddings of each associated mention. Following the methodology of Jia et al. (2019), a soft version of the max function is utilized to compute the embeddings of entities:
where is an entity comprising multiple mentions, is the number of mentions for entity .
@@ 4.2.4 Collaborative Prediction Layer
where is the row vector composed of all , is the total count of sentences contained within the document. and are mainly used in conjunction with Kullback Leibler (KL) divergence for loss calculation.